x^2 -y^2=a^2 ------------> rectangular hyperbola xy=c^2 ----------> rectangular hyperbola where c^2=( a^2 ) / 2--------->How ????????

x^2 -y^2=a^2  ------------> rectangular hyperbola
 
xy=c^2  ----------> rectangular hyperbola
 
where c^2=( a^2 ) / 2--------->How ????????

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
when axes are rotated through 45 degree in clock wise direction.
hence (x -y)(x+y) = a^2
 
(x-y) /sqrt2 (x + y)/ sqrt2 = a^2/2
 
Now XY = c^2
hence on comparison c^2 = a^2/ 2